OpenStudy (anonymous):

If (x-a) is a factor of (x^3- 5x^2- 2x+ 24), what number could be a?

OpenStudy (anonymous):

find the root of x^3- 5x^2- 2x+ 24

OpenStudy (anonymous):

find the roots of this equation.all the roots found can be a

OpenStudy (amoodarya):

x^3- 5x^2- 2x+ 24=(x+2)(x-3)(x-4) x^3- 5x^2- 2x+ 24=(x-a)g(x) +r --->r=0 so a= -2, or +3, or+4